five freinds drove 60 miles per hour to California. on the way home they took the same route but drove 75 miles per hour. how many mile did they drive to California if the round trip took 10 hrs.
Found 2 solutions by jorel1380, stanbon:
Answer by jorel1380(2518)

n/60+n/75=10
5n+4n=3000
9n=3000
n=3000/9=333.33 miles
333.33/60+333.33/75=5.56+4.44=10..

Answer by stanbon(57361)

five friends drove 60 miles per hour to California. on the way home they took the same route but drove 75 miles per hour. how many miles did they drive to California if the round trip took 10 hrs.
------------------------------
To DATA:
distance = x miles ; rate = 60 mph ; time = d/r = x/60 hrs.
--------------------
From DATA:
distance = x miles ; rate = 75 mph ; time = d/r = x/75 hrs.
----------------
Equation:
x/60 + x/75 = 10 hrs
---
75x + 60x = 10*60*75
---
x = 333 1/3 miles (distance to CA)
===================================
Cheers,
Stan H.
============